OpenStudy (anonymous):

Solve 10x + 75 = 150. Round your answer to the nearest hundredth

OpenStudy (anonymous):

Isolate the variable x. 1. Subtract 75 from both sides 2. Divide by 10 on both sides

OpenStudy (goformit100):

@PeterPan

OpenStudy (anonymous):

so easy. 10x + 75 = 150 -75 from 75 & 150 150 - 75 = 75 10x = 75 Divide 10x from both sides 7.5 = x

OpenStudy (anonymous):

Duhhhhh. It's easy as making cereal :P

OpenStudy (goformit100):

just solve for x